___Fe2O3(s) + ___CO(g) à ___Fe(s) + ___CO2(g) 

How many grams of iron metal can be produced from the reaction of 10.0 grams of iron (III) oxide and carbon monoxide? 

How many grams of iron metal can be produced from the reaction of 10.0 grams of carbon monoxide with iron (III) oxide? 

Compare the amount of iron produced in parts a and b. Which is less? 

What is the limiting reactant? 

What is the excess reactant? 

What is the theoretical yield of iron in this reaction? 

How many grams of limiting reactant are unused when the reaction is complete? 

How many grams of the excess reactant are unused when the reaction is complete? 


1
Expert's answer
2021-01-26T03:23:13-0500

Fe2O3(s) +3CO(g) --------- 2Fe(s) + 3CO2(g).


a)

Given moles of Fe2O3 = 10 / 160 = 1 / 6 mol.

So moles of Fe = ( 2 / 6 ) mol . = 0.333 mole . Answer .


b ) 

Given .Moles of CO = 10 * / 28 Mol .

So moles of Fe = ( 2 *10 / 28 * 3 ) mol . = 0.238 mole .Answer .


3)

part ( b ) . have smaller product of Fe .


4 )

Fe2O3 is Limiting Reagent .

CO Is Excess Reagent .

No any amount left after reaction of limiting reagent .

Left amount of Excess Reagent = ( 10 / 28 - 1/2 ) mol.